In a family, a couple has a son and a daughter. The age of the father is four times that of his daughter, and the age of the son is half that of his mother. The wife is 8 years younger than her husband, and the brother is seven years older than his sister. What is the age of the mother?
Answer by TimothyLamb(4379)

x = age of dad
y = age of mom
z = age of son
u = age of daughter
---
x = 4u
y = 2z
y = x - 8
z = u + 7
---
put the system of linear equations into standard form
---
x - 4u = 0
y - 2z = 0
x - y = 8
z - u = 7
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = age of dad = 44
y = age of mom = 36
z = age of son = 18
u = age of daughter = 11
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---